### v1

# Quantifying Phase Unbalance and Coordination Impacts on Distribution Network Flexibility

Abstract
The increasing integration of distributed energy resources (DER) provides distribution system operators (DSO) with new flexible resources to support more efficient operation and planning of distribution networks. To utilise these resources, various DER flexibility aggregation methods have been proposed in the literature, such as aggregated P-Q flexibility areas at the interface with other networks. However, whereas focusing on estimating the limits of flexibility services, existing studies make the critical assumption that all available flexible units are perfectly coordinated to jointly provide flexibility and manage network constraints. Moreover, due to the extensive use of single-phase power flow analysis, the impacts of phase unbalance on DER flexibility aggregation remain largely unexplored. To address these gaps in knowledge, this work proposes a framework for modelling flexibility services in low voltage (LV) distribution networks which enables explicitly imposing voltage unbalance and phase coordination constraints. The simulations, performed for an illustrative 5-bus system and a real 221-bus LV network in the UK, demonstrate that a significant share (over 30%) of total aggregated DER flexibility potential may be unavailable due to voltage unbalances and lack of coordination between DER connected to different phases.
Index Terms:
Aggregated flexibility, distribution network, distributed energy resources (DER), flexibility services, low voltage networks, phase unbalance, voltage unbalance.
I Introduction
Modern distribution systems are developing with the rapid integration of distributed energy resources (DER), especially in low voltage (LV) networks [1, 2, 3]. Controllable DER, such as battery energy storage systems, prosumers, electric vehicle aggregators, etc., have the technical ability to adjust their power exchange with the grid and, therefore, can provide flexibility services, e.g., by balancing electricity supply and demand [4, 5]. These services can be used to improve the planning and operation of distribution systems, as well as be aggregated and traded between distribution and transmission system operators (DSO and TSO) [6, 7, 8].
Flexibility aggregation models are considered key for providing both local flexibility services (e.g., within LV feeders) and system-level services (e.g., at the primary substations or TSO-DSO interfaces) [9, 10, 11, 12, 13, 3, 14, 5, 15, 16, 17, 18, 19, 20]. The idea behind such models is to estimate the limits of flexible active and reactive power exchanges at certain locations of a distribution network. By estimating these limits, DSO can clear flexibility markets while managing network constraints. The sets of all feasible power exchanges (operating points) in the P-Q space are known as aggregated P-Q flexibility areas or nodal operating envelopes [5, 14].
Despite the advances in DER flexibility aggregation models, there are several critical assumptions made in the literature regarding DER operation and coordination in LV networks. Specifically, this work focuses on the following two problems:
  • Problem #1: Existing studies on DER flexibility aggregation and P-Q flexibility areas calculation assume that all flexibility providers (referred to as flexible units in this work) can be perfectly coordinated [9, 10, 11, 12, 5, 15, 16, 17, 18, 19, 20]. This assumption stems from the use of centralised aggregation methods, where units at different locations can jointly adjust their power generation and consumption in order to manage voltage constraints and keep the network’s operation feasible. However, in real distribution systems, flexible units connected to different phases across the network may not be perfectly coordinated due to low observability, data exchange issues, single-phase connection, behind-the-meter devices, and other constraints, as discussed in [14, 13]. Therefore, there is a problem of incorrect estimation of flexible capacity in distribution networks due to insufficient DER coordination, which can be exacerbated by the next problem.
  • Problem #2: There is a lack of understanding of potential phase unbalance impacts on flexibility services provision. This is partly due to the fact that many existing studies on TSO-DSO flexibility markets rely on single-phase power flow models to simulate flexibility in distribution systems, thus assuming that they are perfectly balanced [9, 10, 11, 12, 5, 14, 15, 16, 17, 18, 19, 20]. Other works that do consider the unbalanced nature of distribution networks, e.g., [13, 3], use three-phase optimal power flow (OPF) models for analysing flexibility in LV distribution networks. Such models are capable of managing network constraints across multiple phases and ensuring the feasibility of DER operation in unbalanced networks. However, they do not specifically assess the impacts of phase unbalance on flexibility services, e.g., the impact of voltage unbalance on aggregated P-Q flexibility limits. Considering the coordination problem between flexible units, phase unbalance can become a major bottleneck in aggregating DER flexibility.
These problems are present in real distribution systems as LV distribution networks are inherently unbalanced and the increase of single-phase uncoordinated distributed generation will further aggravate existing voltage unbalances [21, 22]. For example, suboptimal selection of the DER connection phase can lead to voltage unbalance and limit DER capacity [23]. The need for simulating DER flexibility using accurate three-phase power flow models is acknowledged by the literature [24, 2, 25, 3, 26]. However, existing studies that exploit three-phase power flow models focus on assessing the benefits of DER flexibility services or characterising operating envelopes (limits) for prosumers. Despite recent work [27], there has been no attempt to analyse the aggregated P-Q flexibility for unbalanced LV distribution networks. Note that although P-Q flexibility areas were calculated using a three-phase OPF model in [27], the impact of voltage unbalance constraints was not explicitly quantified. Moreover, no studies have investigated the loss of coordination between DER connected to different phases and its consequences for voltage unbalance and the feasibility of flexibility services provision.
In this regard, this paper investigates the provision of flexibility services in unbalanced LV distribution networks and proposes a framework for quantifying the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. At the core of the framework lies an accurate nonlinear three-phase AC OPF model designed for simulating the operation of flexible units connected to different phases. To characterise feasible flexibility services available in distribution networks, the concept of aggregated P-Q flexibility areas is applied. Then, voltage unbalance constraints are explicitly included to limit the deviations in phase voltages caused by the provision of flexibility services. Finally, phase coordination constraints are imposed to simulate situations where flexible units connected to different phases cannot perform coordinated power adjustment to maximise flexibility services and keep voltage unbalance at acceptable levels. The combination of these models and concepts allows to directly translate voltage unbalance and phase coordination constraints as reductions of the aggregated P-Q flexibility.
The models supporting the framework have been developed in Julia programming language and published as an open-source tool. The tool, named 3FlexAnalyser.jl, and the case studies are available at [28].
To the best of the authors’ knowledge, this work is the first detailed analysis of phase unbalance and DER coordination impacts on aggregated P-Q flexibility in LV distribution networks. Specifically, the paper makes the following contributions:
  • Contribution #1: A novel framework is proposed for quantifying the impacts of phase unbalance and DER coordination on aggregated flexibility in LV distribution networks. The framework comprises three-phase AC power flow models, the concept of P-Q flexibility areas, and voltage unbalance constraints. The combination of these models and concepts allows to directly translate voltage unbalance and phase coordination constraints as reductions in the aggregated P-Q flexibility.
  • Contribution #2: New quantification analysis is performed to demonstrate that phase unbalance and lack of coordination between DER connected to different phases can lead to significant losses in potential aggregated DER flexibility. Simulations performed for an illustrative 5-bus system and a real 221-bus LV network in the UK show that over 30% of total aggregated DER flexibility potential may be unavailable.
Moreover, through extensive simulations for case studies with different combinations of voltage unbalance and phase coordination constraints, the following findings have been formulated:
  • Finding #1: Flexibility services in LV distribution networks can be significantly constrained due to inherent load unbalances, even in cases where flexible resources provide balanced output.
  • Finding #2: Several small single-phase flexible units have the potential to provide more flexible power than a few larger three-phase balanced flexible units. However, perfect coordination of single-phase units is required to achieve this potential. Without coordination, single-phase flexible units can exacerbate existing voltage unbalances.
  • Finding #3: The worst conditions for providing flexibility services include the lack of coordination between flexible units connected to different phases and tight voltage unbalance constraints. In such cases, flexible units cannot effectively manage voltage unbalance across different phases and locations, which results in the infeasibility of services provision.
The rest of the paper is organised as follows. Section II introduces the modelling framework for simulating DER operation in LV unbalanced networks and quantifying the impacts of voltage unbalance and phase coordination constraints. In Section III, extensive simulations are performed for an illustrative 5-bus distribution system and a real 221-bus LV network in the UK. Section IV further discusses the aspects of flexibility provision in unbalanced LV distribution networks. Section V concludes the paper.
II Modelling Framework
This section introduces the model formulation and modelling assumptions used to simulate flexibility services in three-phase unbalanced LV distribution networks. The framework consists of three main components: 1) first, an accurate nonlinear AC optimal power flow (OPF) model is selected to represent distribution networks with DER and flexible resources, 2) then, the concept of P-Q flexibility areas is adapted to quantify flexibility services available in the networks, 3) finally, voltage unbalance and phase coordination constraints are imposed to analyse their impacts on flexibility services.
To accurately simulate LV distribution networks with flexible units connected to different phases, a multi-phase flexibility estimation model (1a)-(1p) is formulated as a nonlinear programming problem (NLP). The formulation is based on the bus injection model widely applied in unbalanced distribution network optimisation studies [29, 30]. Variables of the model are defined for each phase $\phi$ and include complex voltages and complex power injections at each bus $i$, $V_{i,\phi}$, $s_{i,\phi}$, complex power flows for each line $(i,j)$, $s_{ij,\phi}$, active and reactive power of generators $P_{g,i,\phi}$, $Q_{g,i,\phi}$, and active and reactive power of flexible units $P_{f,i,\phi}$, $Q_{f,i,\phi}$. Voltage unbalances for the selected set of buses $\mathcal{N}^{\text{vu}}$ are explicitly modelled using the positive and negative voltage sequences, ${V_{1}}_{i},{V_{2}}_{i}$, and voltage unbalance factors (VUF), $VUF_{i}$. Note that the original mathematical formulation of the three-phase power flow problem is adapted from the package PowerModelsDistribution.jl [30]. It is then extended by including flexible units, voltage unbalance limits, and phase coordination constraints.
The objective function (1a) minimises or maximises the network’s active and reactive power injections at the reference bus, where coefficients $\alpha^{p}_{i}$ and $\alpha^{q}_{i}$ are introduced to control the optimisation directions in the P-Q space. In this formulation, reference bus $i^{\text{ref}}$ is the location for which flexibility from multiple flexible units is aggregated. Thus, the objective function with adjustable coefficients allows to iteratively (point by point) estimate the limits of aggregated flexibility in the P-Q space at the reference bus [14, 31]. Equations (1b)-(1k) represent the nonlinear AC OPF constraints. Specifically, voltage limits are imposed in (1b), complex power flows and bus injections are defined in (1c), (1d), where $Y_{ij}$ is the admittance matrix. Active and reactive power balance constraints for each node are introduced in (1e), (1f), where $d\in\mathcal{D}$ represents loads of the network. Line capacity limits are included in (1g). Active and reactive power of generators and flexible units is limited in (1h), (1i) and (1j), (1k). The formulated AC OPF model (1a)-(1k) enables estimating the limits of flexibility services for the selected phase and bus of a distribution network, subject to available flexible resources and network constraints.
[…]
Variables: (for each phase $\phi\in\{a,b,c\}$)
$V_{i,\phi}$ $i\in\mathcal{N}$
$s_{i,\phi}$ $i\in\mathcal{N}$
$s_{ij,\phi}$ $(i,j)\in\mathcal{L}$
$P_{g,i,\phi},Q_{g,i,\phi}$ $i\in\mathcal{N},g\in\mathcal{G}\quad$
$P_{f,i,\phi},Q_{f,i,\phi}$ $i\in\mathcal{N},f\in\mathcal{F}$
${V_{1}}_{i},{V_{2}}_{i},{VUF}_{i}$ $i\in\mathcal{N}^{\text{vu}}$
Objective:
$\min\enskip\alpha^{p}_{i}\Re(s_{i,\phi})+\alpha^{q}_{i}\Im(s_{i,\phi})\quad$ $i=i^{\text{ref}}\quad\vspace{1\jot}$ (1a)
AC OPF constraints:
$V^{\min}\leq\left|V_{i,\phi}\right|\leq V^{\max}$ $\forall i\in\mathcal{N}$ (1b)
$s_{ij,\phi}={diag\left[V_{i,\phi}\left(V_{i,\phi}-V_{j,\phi}\right)^{H}{Y_{ij}}^{H}\right]}\quad$ $\forall(i,j)\in\mathcal{L}\quad$ (1c)
$s_{i,\phi}=\sum_{(i,j)\in\mathcal{L}}s_{ij,\phi}$ $\forall i\in\mathcal{N}\quad$ (1d)
$\sum_{g,d,f}^{\mathcal{G},\mathcal{D},\mathcal{F}}(P_{g,i,\phi}-P_{d,i,\phi}+P_{f,i,\phi})=\Re(s_{i,\phi})$ $\forall i\in\mathcal{N}$ (1e)
$\sum_{g,d,f}^{\mathcal{G},\mathcal{D},\mathcal{F}}(Q_{g,i,\phi}-Q_{d,i,\phi}+Q_{f,i,\phi})=\Im(s_{i,\phi})\enskip\enskip$ $\forall i\in\mathcal{N}$ (1f)
$\left|s_{ij,\phi}\right|\leq s_{ij,\phi}^{\max}$ $\forall(i,j)\in\mathcal{L}$ (1g)
$P_{g,i,\phi}^{\min}\leq P_{g,i,\phi}\leq P_{g,i,\phi}^{\max}$ $\forall g\in\mathcal{G}$ (1h)
$Q_{g,i,\phi}^{\min}\leq Q_{g,i,\phi}\leq Q_{g,i,\phi}^{\max}$ $\forall g\in\mathcal{G}$ (1i)
$P_{f,i,\phi}^{\min}\leq P_{f,i,\phi}\leq P_{f,i,\phi}^{\max}$ $\forall f\in\mathcal{F}$ (1j)
$Q_{f,i,\phi}^{\min}\leq Q_{f,i,\phi}\leq Q_{f,i,\phi}^{\max}$ $\forall f\in\mathcal{F}\vspace{1\jot}$ (1k)
Voltage unbalance limits:
${V_{1}}_{i}=\frac{{V_{i,\{a\}}}+\boldsymbol{a}{V_{i,\{b\}}}+\boldsymbol{a}^{2}{V_{i,\{c\}}}}{3}$ $i\in\mathcal{N}^{\text{vu}}$ (1l)
${V_{2}}_{i}=\frac{{V_{i,\{a\}}}+\boldsymbol{a}^{2}{V_{i,\{b\}}}+\boldsymbol{a}{V_{i,\{c\}}}}{3}$ $i\in\mathcal{N}^{\text{vu}}$ (1m)
$VUF_{i}=\frac{{V_{2}}_{i}}{{V_{1}}_{i}}\times 100\%$ $i\in\mathcal{N}^{\text{vu}}$ (1n)
$VUF_{i}\leq\overline{VUF}$ $i\in\mathcal{N}^{\text{vu}}\vspace{1\jot}$ (1o)
Phase coordination constraints:
$\sum_{i\in\mathcal{N},f\in\mathcal{F}}P_{f,i,\varphi}=0,\enskip\sum_{i\in\mathcal{N},f\in\mathcal{F}}Q_{f,i,\varphi}=0$ $\forall\varphi\neq\phi\vspace{-1\jot}$ (1p)
[…]
To characterise the flexibility services provided by LV distribution networks, the concept of P-Q flexibility areas is applied to the context of multi-phase flexibility provision. Mathematically, the operation limits of multiple flexible units can be described using the Minkowski addition of their P-Q capability sets [5, 36]. That is, the limits of the aggregated DER flexibility in the P-Q space provided at the reference bus can be estimated as the Minkowski addition of individual flexibility limits of all available flexible units. However, the Minkowski addition does not consider network constraints. To accurately estimate the aggregated DER flexibility considering network constraints, the formulated optimisation model (1a)-(1p) can be solved iteratively, and the boundary of the network flexibility area at the reference bus can be estimated with the desired level of granularity. Multiple algorithms have been proposed for constructing the boundaries of flexibility areas, including Monte Carlo simulations and various optimisation methods [37, 18, 19]. In this work, the $\epsilon$-constraint method is chosen as a simple and straightforward algorithm for approximating the P-Q flexibility area by considering $k$ intervals [16, 14, 31]. At each interval, the feasible P-Q space of the aggregated DER flexibility at the reference bus $i^{\text{ref}}$ is constrained by inequalities (2a) or (2b).
$p_{i,\phi}^{k}-\epsilon\leq\Re(s_{i,\phi})\leq p_{i,\phi}^{k}+\epsilon\qquad$ $i=i^{\text{ref}},\phi=\phi^{\text{ref}}\quad$ (2a)
$q_{i,\phi}^{k}-\epsilon\leq\Im(s_{i,\phi})\leq q_{i,\phi}^{k}+\epsilon\qquad$ $i=i^{\text{ref}},\phi=\phi^{\text{ref}}\quad$ (2b)
These inequalities fix P or Q components of the aggregated flexibility (with a negligibly small margin $\epsilon$) to specific values of interval $k$, $p_{i,\phi}^{k}$ or $q_{i,\phi}^{k}$. Then, model (1a)-(1p) can be solved twice to find the maximum and minimum of the non-fixed flexibility component for this interval. By iteratively imposing the $\epsilon$ constraints and solving model (1a)-(1p), the boundary of the P-Q flexibility area can be calculated as a piece-wise linear approximation with $2k$ operating points. This method gives an easy control over the accuracy (granularity) of the aggregated flexibility estimation. With larger $k$, the accuracy of P-Q flexibility area estimation increases, but the computational cost also grows as the higher numbers of AC OPF problems have to be solved.
Figure 1: Flowchart of the proposed multi-phase flexibility quantification framework: flexibility services from DER are estimated using the concept of P-Q flexibility areas; then, the impact of phase unbalance and DER coordination is analysed by introducing the corresponding constraints.
Note that existing studies on TSO-DSO flexibility markets estimate the P-Q flexibility areas under the assumption of balanced power flow. In this work, the concept of flexibility areas is extended to a multi-phase context: flexibility service limits are estimated in the P-Q space individually for each phase. This requires selecting a reference phase $\phi^{\text{ref}}$ for which the limits of P-Q flexibility will be calculated. The multi-phase formulation enables quantifying the impacts of voltage unbalance and phase coordination constraints. A high-level overview of the proposed framework is presented in Fig. 1. The inputs include network data, available flexible resources, load and flexibility unbalances, VUF limit, and phase coordination assumptions. Then, the limits of flexibility services from DER are estimated using the concept of P-Q flexibility areas. Next, voltage unbalance and phase coordination constraints are imposed and the P-Q flexibility areas are estimated again. Finally, the aggregated flexibility is compared for cases with and without these constraints and the impacts of phase unbalance and DER coordination are quantified (as reductions in the aggregated P-Q flexibility areas, e.g., in kVAr2).
III Case Studies
In this section, the proposed framework is demonstrated using two case studies: (i) a simple 5-bus system with DER flexibility available at a single location; (ii) a real 221-bus distribution network in the UK with 12 flexible units. Data for these cases is available at [38] and is also included in the code repository [28]. For each system, the limits of flexibility services are characterised via the concept of aggregated P-Q flexibility areas. Then, the impacts of voltage unbalance and phase coordination are quantified and the main findings are discussed. All simulations have been performed with JuMP 1.11.0 for Julia 1.6.1 language and Ipopt 3.14.4 solver.
III-A Case Study: Illustrative 5-bus System
To illustrate the principles of flexibility services estimation considering phase unbalance, a simple network with 5 buses is constructed, visualised in Fig. 2. The network has four loads and one flexibility provider. Each load has connections to three phases, which enables to simulate balanced network operation or to create phase unbalance due to uneven allocation of single-phase loads. [^2] In this work, it is assumed that a three-phase balanced voltage source is connected to the primary substation, representing power supply from a medium voltage or high voltage network. That is, the voltage of the source is set to 1.0 p.u. for each phase. The inclusion of unbalanced power supply sources is left for future research. The flexible unit also has connections to all phases and therefore can be used to simulate balanced or unbalanced flexible power output.
Specifically, the following network conditions are considered in simulations for this case study:
  • Balanced network: All loads are equal to 5 kW and are perfectly balanced, that is, $P_{d,i,\phi}=5.0$ $\forall i\in\mathcal{N},\phi\in\{a,b,c\}$. The load power factor is assumed to be constant and equals 0.91.
  • Unbalanced network: Load L2 has different active power consumption per phase, {7.0, 4.5, 3.5} kW, with a power factor of 0.91. All other loads are equal to 5 kW and are perfectly balanced.
  • Balanced flexibility: The flexible unit is a three-phase device producing the same flexible power output per each phase, with limits ±8 kW and ±8 kVAr, that is, $P_{f,\phi}\in[-8.0,8.0],$ $Q_{f,\phi}\in[-8.0,8.0]$ $\forall\phi\in\{a,b,c\},$ $P_{f,\{a\}}=P_{f,\{b\}}=P_{f,\{c\}},$ $Q_{f,\{a\}}=Q_{f,\{b\}}=Q_{f,\{c\}}$.
  • Unbalanced flexibility: Three independent single-phase flexible units are connected at bus 3. These units can have different active and reactive power outputs (with limits ±8 kW and ±8 kVAr), thus providing unbalanced flexibility.
Figure 2: Case study: illustrative 5-bus system with four loads and one flexible unit. Each load can consume different power across the phases, thus creating phase unbalance. The flexible unit is assumed to produce a balanced output (to simulate a three-phase flexible device) or an unbalanced output (to simulate three independent flexibility providers connected to different phases).
First, the impact of load unbalance on flexibility services is analysed. The flexible unit is assumed to be a three-phase device producing the same flexible power output per phase. That is, by design, this unit can provide only balanced flexibility services. The question is, does the load unbalance, inherent in LV distribution networks, affect such services? To address this question, the aggregated P-Q flexibility area of the network was calculated for cases with balanced and unbalanced loads, as shown in Fig. 3. Each boundary was approximated by 80 points, i.e., the flexibility estimation model (1a)-(1p) was iteratively solved 80 times. In the case of a balanced network, the flexibility service is less constrained and has the largest P-Q area (displayed in grey colour). [^3] Note that for a perfectly balanced network, there is no need to display flexibility areas per each phase as these areas coincide. In the case of an unbalanced network, the operation of the flexible unit is constrained differently in each phase. Therefore, the limits of the flexibility service are reduced, and the P-Q areas become different in the phases (as highlighted in blue, orange, and green colours). The quantification of the flexibility reduction due to load unbalance is given in Table I, where the areas of aggregated flexibility are calculated in kVAr2. The results demonstrate that the areas are significantly reduced due to the introduced load unbalance (36-38% reduction for the 5-bus system). It follows that balanced flexibility services provided by three-phase devices can be constrained by load unbalance present in distribution networks.
Figure 3: Analysis of the load unbalance impact on flexibility services provided by the three-phase balanced flexible unit in the 5-bus system. The services are characterised by the aggregated network flexibility area in the P-Q space at the primary substation. The cross markers correspond to the initial operating point (with no flexibility provision), while the coordinates represent the total network’s power consumption.
TABLE I: Load Unbalance Impacts on Aggregated P-Q Flexibility Provided by a Balanced Three-phase Unit in the 5-bus System
Simulations for balanced flexibility
Aggregated P-Q flexibility area,
kVAr2 (%)
Balanced network, any phase 267.96 (100.00%)
Unbalanced network, $\phi=\{a\}$ 177.03 (66.07%)
Unbalanced network, $\phi=\{b\}$ 165.86 (61.90%)
Unbalanced network, $\phi=\{c\}$ 166.16 (62.01%)
Second, the impact of voltage unbalance limits on flexibility services is analysed. An unbalanced network with three single-phase flexible units is considered. That is, both loads cause phase unbalance and the units can provide unbalanced flexibility services. Then, the VUF constraints (1l)-(1o) are imposed to control voltage unbalance caused by the uneven allocation of single-phase loads and unbalanced output of single-phase flexible units. The aggregated P-Q flexibility areas for VUF in the range 0.1-1.5% are shown in Fig. 4.444Note that VUF limits below 1% are unrealistically low – no grid codes regulate such low voltage unbalances. In this work, low VUF limits are included in the analysis for the sake of illustrating the principles of flexibility provision under phase unbalance. [^5] In the 5-bus system, VUF constraints (1l)-(1o) are imposed on each bus. For simplicity, only phase $\phi=\{b\}$ is displayed in the figure – flexibility services in other phases have similar dependencies on VUF constraints.
Note that the area displayed in Fig. 4 is larger than the areas for balanced flexibility services analysed in Fig. 3 and Table I. The three single-phase flexible units can provide more flexibility (with an area of 306.04 kVAr2) than a single three-phase balanced unit of the same capacity. It follows that, if coordinated, multiple single-phase units can outperform three-phase units in maximising the limits of aggregated flexibility. It is also worth noting that VUF limits do not have a major impact on the flexibility provided by multiple single-phase units. Even under the extreme unrealistic VUF limit of 0.1%, the P-Q flexibility area (highlighted in red) decreases by only 6.91%. This happens due to the assumption of perfect unit coordination. Under this assumption, all units jointly manage voltage constraints and maximise flexibility services: any unbalance constraining flexibility in one phase can be mitigated by coordinated actions of units in other phases.
Figure 4: Analysis of the voltage unbalance impact on flexibility services in the 5-bus system, for phase $\phi=\{b\}$. The cross marker corresponds to the initial operating point (with no flexibility provision), while the coordinates represent the total network’s power consumption. The services are provided by three single-phase flexible units connected to different phases. All units are assumed to be perfectly coordinated.
However, for current real distribution systems, the perfect coordination assumption may be unreasonable since 1) single-phase flexible units can be connected at different locations and belong to different owners and 2) required automation and communication technologies are not widely adopted. To analyse the lack of unit coordination, phase coordination constraint (1p) is imposed. Under this constraint, flexibility service in phase $\phi$ can be provided only by the single-phase unit connected to this phase. Other units cannot manage the voltage unbalance or assist in maximising network flexibility. The impact of phase coordination on P-Q flexibility areas is shown in Fig. 5, where VUF constraints (1o) in the range 0.1-1.0% are iteratively imposed. The findings of this simulation are twofold: (i) without coordination between units connected to different phases, the limits of aggregated flexibility reduce significantly, for example, 60.76% flexibility area reduction for phase $\phi=\{b\}$; (ii) voltage unbalance limits have a major impact on flexibility services provided by units with a lack of coordination. It follows that a lack of phase coordination combined with voltage unbalance limits creates the worst conditions for providing flexibility services: single-phase units cannot effectively manage voltage unbalance across the network, which results in the infeasibility of services provision.
Figure 5: Analysis of the phase coordination and voltage unbalance impacts on flexibility services in the 5-bus system: (a) for phase $\phi=\{a\}$; (b) for phase $\phi=\{b\}$; (c) for phase $\phi=\{c\}$. The cross markers correspond to the initial operating point (with no flexibility provision), while the coordinates represent the total network’s power consumption. The services are provided by three single-phase flexible units. Phase coordination constraints imposed do not allow units connected to different phases to jointly provide flexibility and manage voltage unbalance, thus reducing flexibility services for each phase.
a)Refer to caption  b)Refer to caption  c)Refer to caption
The findings illustrated in this subsection emphasise the need for considering phase unbalance impacts on flexibility services in distribution networks. If ignoring these impacts, the potential of DER flexibility services may be greatly overestimated. The following subsection presents a realistic case study to demonstrate the scalability of the proposed framework.
III-B Case Study: 221-bus LV Distribution Network in the UK
The proposed flexibility quantification framework is applied to a real anonymised 0.4 kV distribution network located in the North West of England [38]. The network is visualised in Fig. 6 as a graph using the force-directed graph layout algorithm ForceAtlas2 [39]. The figure illustrates the network topology and the location of power demands across the network, where circles of different sizes correspond to the total apparent power consumed at each bus. The smallest circles indicate buses with no loads, while the largest circles stand for buses with a load of 3.52 kVA. The distribution of loads across phases is given in Table II. The network is supplied from a single three-phase balanced source, the primary substation (displayed as a triangle in Fig. 6). Yet, due to the uneven distribution of single-phase loads, the network experiences a voltage unbalance, with the highest VUF of 0.78% over all buses. Such a low level of voltage unbalance is considered acceptable according to existing grid codes [34, 35]. However, this voltage unbalance can still impact the network’s flexibility and limit the operation of DER.
Figure 6: Case study: real 221-bus LV distribution network in the UK. The system has 12 single-phase flexible units (4 units per phase) as highlighted by the corresponding colours.
TABLE II: Unbalanced Load Data for the 221-bus LV Network
Total $\phi_{a}$ $\phi_{b}$ $\phi_{c}$
P, kW 82.00 37.00 24.00 21.00
Q, kVAr 64.55 31.41 17.06 16.08
To investigate the impact of phase unbalance on the network’s flexibility services, 12 single-phase flexible units (four units per phase) were placed in the network as shown in Fig. 6. Each unit is assumed to have a P-Q capability of ±5 kW and ±5 kVAr. That is, units can independently regulate their active and reactive power output in the range of [-5.0,5.0] kW and [-5.0,5.0] kVAr. Then, the aggregated P-Q flexibility areas were calculated per each phase by iteratively solving the flexibility estimation model (1a)-(1p). The impact of phase unbalance and coordination on P-Q flexibility areas is shown in Fig. 7, where VUF constraints (1o) in the range 0.5-1.0% are imposed. [^6] To avoid scalability issues, in the 221-bus system, VUF constraints (1l)-(1o) are imposed only on the 7 buses with the highest voltage unbalances.
Similar to the previous analysis for the simple 5-bus system, the P-Q areas in Fig. 7 indicate that the network’s flexibility is maximal when all single-phase flexible units are perfectly coordinated (as shown by the dashed lines). However, the absence of coordination between units connected to different phases leads to significant losses in the aggregated DER flexibility potential. For example, the P-Q flexibility area for phase $\phi=\{a\}$ reduces by 30% (from 1721.7 to 1211.6 kVA2). Moreover, imposing voltage unbalance constraints (1o) further limits the operation of flexible units and significantly reduces the network’s aggregated flexibility.
Figure 7: Analysis of the phase coordination and voltage unbalance impacts on flexibility services in the real 221-bus system: (a) for phase $\phi=\{a\}$; (b) for phase $\phi=\{b\}$; (c) for phase $\phi=\{c\}$. The cross markers correspond to the initial operating point (with no flexibility provision), while the coordinates represent the total network’s power consumption. The services are provided by 12 single-phase flexible units. Phase coordination constraints imposed do not allow units connected to different phases to jointly provide flexibility and manage voltage unbalance, thus reducing flexibility services for each phase.
a)Refer to caption  b)Refer to caption  c)Refer to caption
Note that in this simulation setting, all flexible units have been assigned the same P-Q capabilities and distributed evenly across phases. Yet, their operation is limited considerably due to phase coordination issues and voltage unbalance. It follows that the provision of flexibility services by DER can be challenging even in LV networks with low load unbalance as flexible units can exacerbate existing voltage unbalances by producing or consuming power. To effectively manage such unbalances and voltage constraints, units connected to different phases must be perfectly coordinated. The simulations performed for the realistic 221-bus network demonstrate that lack of coordination between units significantly reduces the potential of DER flexibility services and can magnify voltage unbalance problems. Other aspects and implications of these findings are discussed in the next section.
IV Discussion
The simulations performed in Section III demonstrate the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. Specifically, it is shown that the total DER flexibility potential can be significantly reduced due to voltage unbalance constraints and lack of coordination between flexible units connected to different phases. This highlights the importance of accurate nonlinear three-phase OPF models for the analysis of the network’s aggregated P-Q flexibility, as such models are capable of simulating voltage unbalances and coordination issues across phases. In this regard, the proposed flexibility quantification framework allows to directly translate voltage unbalance and phase coordination constraints as reductions of the aggregated P-Q flexibility. Thus, it can be used by DSO to realistically predict flexibility available in distribution networks and avoid severe voltage unbalances.
Multiple aspects of flexibility provision in unbalanced LV distribution networks have to be further investigated. One of the key factors affecting the aggregated P-Q flexibility is the relationship between the placement of flexible units and locations of voltage unbalances. For instance, in the 221-bus case study displayed in Fig. 6, the highest load unbalance is happening at the feeder with the flexible unit F5{a} connected. Six more flexible units are placed in this part of the network. Therefore, the design of this experiment inherently creates constraints on the operation of these flexible units due to the nearby voltage unbalances. That is, units must be coordinated when consuming or producing power to manage voltage constraints and keep the network’s operation feasible. For example, units connected to phase $\phi=\{a\}$ (the phase with the highest load) must not increase their power consumption while units in other phases consume power. Otherwise, voltage unbalance in some parts of the network may increase to unacceptable levels or even the voltage limits (1b) for phase $\{a\}$ may be violated. Note that this would not happen if flexible units are located far from the load unbalance, say, at the primary substation. Therefore, the question arises: how does the initial network unbalance impact the network’s P-Q flexibility and how does it restrict flexible resources at different locations? This question has implications for DER planning models. Depending on DSO metrics and objectives, some DER should be placed closer to the primary substation, while other flexible resources should be located closer to existing voltage unbalances to manage voltage constraints there. A similar problem was considered in [23], where it was demonstrated that suboptimal selection of the DER connection phase can lead to voltage unbalance and limit DER capacity.
Regarding the computational aspects of flexibility quantification, the main computational burden comes from the need to iteratively solve the nonlinear optimisation problem (1a)-(1p). A single OPF problem with no voltage unbalance constraints converges in 0.1-0.2 seconds and 5-50 seconds for the 5-bus and 221-bus cases, respectively. This means that the boundary of an aggregated P-Q flexibility area with 50 points can be estimated in less than 2500 seconds (43 minutes). On average, it takes 16 minutes to calculate a single boundary presented in Fig. 7. Note that these values are reported for the simulations performed on a laptop with an Intel Core i7-10510U CPU 1.80GHz and 16 GB of RAM. The flexibility estimation process can be greatly sped up by using more powerful computers and parallel computing (solving OPF problems for different points on different machines). Therefore, DSO can perform this flexibility assessment close to real time.
There also exist computational challenges related to the number of flexible units available in the network and the number of buses on which VUF constraints (1l)-(1o) are imposed. In the 221-bus case study, VUF constraints were imposed only on 7 buses (corresponding to buses with the highest voltage unbalances). The inclusion of VUF limits for all 221 buses dramatically increases the computation time and makes the problem intractable. This paper aims to illustrate the impacts of phase unbalances on flexibility services in distribution networks. Thus, scalability challenges in large networks with numerous DER and load unbalances are left for future research.
V Conclusion
This paper investigates the provision of flexibility services in unbalanced LV distribution networks and proposes a framework for quantifying the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. At the core of the framework lies an accurate nonlinear three-phase AC OPF model designed for simulating the operation of flexible units connected to different phases. To characterise feasible flexibility services available in networks, the concept of aggregated P-Q flexibility areas is applied. The combination of these models and concepts allows to directly translate voltage unbalance and phase coordination constraints as reductions of the aggregated P-Q flexibility. The simulations, performed for an illustrative 5-bus distribution system and a real 221-bus LV network in the UK, demonstrate that a significant share (over 30%) of total aggregated DER flexibility potential may be unavailable due to voltage unbalances and lack of coordination between DER connected to different phases.
A detailed analysis of these simulations enables to formulate several findings explaining the operation of DER in unbalanced distribution networks and their role in aggregated flexibility services. For example, it is found that flexibility services in LV distribution networks can be significantly constrained due to inherent load unbalances, even in cases where flexible resources provide balanced output. Moreover, in cases where flexible units are connected to different phases, they can exacerbate existing voltage unbalances by producing or consuming power. To effectively manage such unbalances and voltage constraints, units connected to different phases must be perfectly coordinated. However, in realistic networks, flexible units cannot effectively coordinate to manage voltage constraints across different locations and phases. Thus, provision of some flexibility services by the network may be infeasible or may lead to severe voltage unbalances.
Future research can further explore different aspects of flexibility aggregation in unbalanced LV distribution networks. For example, the location of flexible units plays a crucial role in managing voltage unbalances. In this regard, it is worth exploring flexibility planning models and metrics for optimising DER placement while capturing the relationship between DER locations and buses with the highest load unbalances. Concerning computational efficiency, the framework can be further improved to better scale with the number of flexibility providers and VUF constraints.